(a) state the null and alternative hypotheses, (b) compute the test

statistic, (c) determine the critical value and sketch the rejection region and non-rejection

region in the normal curve.


1. The cahier of a fast-food restaurant claims that the average amount spent by customers for dinner is

Php 120.00. A sample of 50 customers over a month was randomly selected and it was found out

that the average amount spent for dinner was Php 122.50. Construct the critical regions using a 0.05

level of significance to conclude that the average amount spent by customers is more than Php

120.00. Assume that the population standard deviation is Php6.50.


1
Expert's answer
2022-05-18T09:38:52-0400

1.

a) The following null and alternative hypotheses need to be tested:

"H_0:\mu=120"

"H_1:\mu>120"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.


b) The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{122.50-120}{6.50/\sqrt{50}}\approx2.7196"



c) Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a right-tailed test is "z_c = 1.6449."

The rejection region for this right-tailed test is "R = \{z:z> 1.6449\}."




d) Since it is observed that "z=2.7196>1.6449=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(z>2.7196)=0.003268," and since "p=0.003268<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than 120, at the "\alpha = 0.05" significance level.